OpenStudy (anonymous):

someone please help with algebra 2 I'm desperate

OpenStudy (anonymous):

Compare the graphs of the inverse variations by comparing asymptotes, similar points, lines of reflections, and quadrants. y=-0.7/x and y=-0.9/x